OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x - 9y =0 -x + 3y = -3 A.1 B. 2 <------ C. infinitely many D. 0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Would you like me to explain why?

OpenStudy (anonymous):

yes please @dancerjeff

OpenStudy (anonymous):

okay... If you use the method of elimination... Step 1) Multiply one of the equations so at least one term cancels out when you add. 

OpenStudy (anonymous):

Step 2) Add the top and bottom equation vertically and combine like terms like so... 3x - 9y = 0 + -3x + 9y = -9 _________________________ 0x + 0y = -9 Simplify knowing that 0 multiplied by anything is still zero. 0 + 0 = -9 0 = -9 As you can see, this statement is false. So, there are no or 0 solutions.

OpenStudy (anonymous):

thank you @dancerjeff

OpenStudy (anonymous):

No prob! ;)